OpenStudy (anonymous):

simplify log base 4 of 1024

OpenStudy (jdoe0001):

\(\large \bf log_{\color{red}{ a}}{\color{blue}{ b}}=y\iff {\color{red}{ a}}^y={\color{blue}{ b}}\qquad thus \\ \quad \\ log_{\color{red}{ 4}}{\color{blue}{ 1024}}=\square\implies {\color{red}{ 4}}^\square ={\color{blue}{ 1024}}\quad thus \quad \square =?\)

OpenStudy (jhannybean):

\[\log_4 1024 = x\]Now you have to ask yourself, 4 to what power will equal 1024? putting that down in a formula, we have \[4^x = 1024\] In order to solve for x, we have to get 1024 to look like \(4^n\)

OpenStudy (anonymous):

ok, thanks, i think i get it now, so im just simplifying this log into not a log?

OpenStudy (jhannybean):

\[1024 = 4 \cdot 4^4 = 4^1 \cdot 4^4 = 4^{4+1} = 4^{5}\] And yes , you're using the properties of logarithmic functions to evaluate x.

OpenStudy (jhannybean):

So I'll start you off and you can finish it! \[4^x = 4^5\] Because the bases are the same, we can simply evaluate their powers.

OpenStudy (anonymous):

oh ok, thanks, i thought it was one of these different properties tht i learned ttoday, not ones i learned yesterday, i just had to look through my old notes, thanks

OpenStudy (anonymous):

wouldnt it be 4^5=1024

OpenStudy (jhannybean):

No problem :) And yes, 4^5 = 1024. So x would equal?

OpenStudy (anonymous):

the exponent 5

OpenStudy (jhannybean):

Yes :D x = 5. Good job!

OpenStudy (anonymous):

thanks so much